I am to recall that, when interest is compounded continuously, the balance in an account after t years is given by
A = Pe^rt,
where P is the initial investment and r is the interest rate.
Maya has deposited $600 in an account that pays 5.64% interest, compounded continuously. How long will it take for her money to double?
Can someone please help with this one? Thank you.

Answer by stanbon(48535)

A = Pe^rt,
where P is the initial investment and r is the interest rate.
Maya has deposited $600 in an account that pays 5.64% interest, compounded continuously. How long will it take for her money to double?
-------------------------
2*600 = 600*e^(0.0564t)
2 = e^(0.0564t)
Take the natural log to get:
0.0564t = ln2
t = 12.2898... years
===================
Cheers,
Stan H.